# Remote Sensing Reveals Adoption of Sustainable Rice Farming Practices Across Punjab, India

### v1

@@ 4 Results
4.1 Dimensional Classification Outperforms Combined Approach
Table 1 presents performance metrics across three classification tasks. The dimensional approach, which separates water management practices along sowing (DSR vs. PTR) and irrigation (AWD vs. CF) dimensions, significantly outperforms the combined three-class approach. When classified along their respective dimensions, both DSR and AWD achieve F1 scores of 78% and 80%, respectively. This result confirms our hypothesis that water management practices are more effectively detected when classified along their natural dimensions rather than being addressed simultaneously. However, as demonstrated in subsection 4.2, AWD classification relies predominantly on lag features, and the models fail to capture irrigation frequency directly, raising concerns about their generalizability. In contrast, DSR classification successfully exploits both lag and sowing features, indicating more robust pattern recognition.
Temporal Ranges Sowing Irrigation F1 (weighted, %)
Row Dates M J J A S O N D Lag (%) DSR PTR CF AWD ALL DSR AWD
1 May 1 - Aug 15 ~50% $\checkmark$ $\checkmark$ $>$ $\times$ 51.88 73.97* 73.48*
2 Jun 1 - Aug 30 ~50% $>$ $\checkmark$ $\checkmark$ $<$ 57.75 71.99 76.31
3 Jun 1 - Sep 15 ~50% $>$ $\checkmark$ $\checkmark$ $<$ 57.75 77.82* 76.31
4 Jun1 - Oct 15 ~60% $>$ $\checkmark$ $\checkmark$ $\checkmark$ 59.62 76.88 76.96
5 July 1 - Oct 15 ~50% $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 55.09 72.84 71.94
6 Aug 1 - Oct 15 ~33% $\times$ $\times$ $\checkmark$ $\checkmark$ 48.94 69.64 70.92
7 Aug 1 - Nov 15 ~50% $\times$ $\times$ $\checkmark$ $\checkmark$ 43.39 72.84 70.69
8 Sep 1 - Dec 15 ~50% $\times$ $\times$ $\checkmark$ $\checkmark$ 49.86 67.68 66.20
9 Oct 1 - Dec 15 ~33% $\times$ $\times$ $<$ $\checkmark$ 51.93 65.80 64.27
10 Aug 1 - Dec 15 ~60% $\times$ $\times$ $\checkmark$ $\checkmark$ 47.00 73.07 68.30
11 Jun 1 - Dec 15 ~33% $>$ $\checkmark$ $\checkmark$ $\checkmark$ 59.62 75.34** 79.12**
12 May 1 - Dec 15 ~90% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 55.29 73.46** 80.33**
Table 2: Effect of temporal range selection on accuracy. Features are generated for different temporal ranges to analyze the effect on outcomes by manipulating the availability of lag and practice features. Cell colors for months represent how much of the month is covered for that row (blank=none, green=full, orange=half). Values in bold indicate the top-2 performers for each classification task. Nominal sampling frequency of 7 days; *= a sampling frequency of 4 days, **=10days.
4.2 Temporal Features Drive Classification Performance
Our analysis reveals two distinct types of temporal features affecting classification performance:
Lag features capture differences in planting schedules between farms. PRANA collected an independent dataset documenting planting dates across practice types (Supp. Fig. S3), which reveals that while substantial variance exists in planting dates across the state, each practice type exhibits a considerably tighter distribution. These temporal offsets provide significant discriminatory power; however, this discriminatory capability depends on the persistence of such patterns over different seasons.
Practice features reflect water management patterns unique to each technique. Sowing features (active for DSR/PTR) are most distinct during the first weeks after planting, while AWD/CF irrigation patterns emerge later in the growing season, as a function of their irrigation frequency, which typically ranges from 4 to over 10 on average (Fig. 4).
To assess the relative importance of these features, we selectively constrained features to different temporal windows during model training (Table 2), thereby highlighting the differential prominence of each feature type. For sowing classification, models exposed to sufficient lag features (>~50%) and containing the majority of DSR and PTR sowing features (rows 1, 3, 10-11) demonstrate strong performance. When DSR features are absent (rows 4-9) or insufficient lag features are available (row 2), DSR classification performance deteriorates markedly. This pattern indicates that sowing models leverage both practice and lag features, conferring partial resilience to changes in cropping calendars. Furthermore, the optimal model appears robust to sampling frequency selection, as the top two performing models employ different sampling frequencies (f=4,7 for rows 3,4).
For irrigation classification, however, we observe strong dependence on lag features, with performance peaking when all features are available (row 11) and declining progressively as lag features are reduced (rows 11 >> row 8). Moreover, performance collapses in the absence of sowing features (rows 5-9), while models maintain reasonable performance even without irrigation features (rows 1-2). This indicates that AWD performance is predominantly driven by lag features and largely disregards irrigation features. This counterintuitive result suggests that AWD irrigation signals are not effectively captured with Sentinel-1A’s 12-day revisit frequency, which proves insufficient for detecting AWD irrigation cycles that typically occur every 4-10 days. This interpretation is further supported by evidence from rows 11-12, where increased sampling frequency to 10 days (i.e. lower temporal resolution incapable of capturing high frequency AWD events) yields optimal model performance.
4.3 Error Analysis Reveals Impact of Temporal Overlap
For both DSR and AWD classification, errors occur predominantly in samples with temporally similar planting dates (Fig. S4). This error pattern confirms that while our dimensional approach effectively addresses most planting date heterogeneity, classification remains challenging when practices exhibit similar temporal signatures. The elevated error rate in AWD classification in the presence of irrigation features [^3] These rows were additionally modeled with a higher sampling frequency of 4 days, yielding even poorer results. (Table 2 row 8-10) further corroborates our finding that current satellite revisit frequencies inadequately capture irrigation cycles.
4.4 Optimizing for Large-Scale Monitoring
For operational monitoring at scale, we tested various satellite data configurations. We found that ascending orbit data consistently outperforms descending orbit data for DSR classification, with no significant benefit from combining both orbits. Additionally, including climatology variables (temperature and precipitation) provides no performance boost (Fig. S5). These findings suggest that a streamlined approach using only ascending orbit data without climatology variables is optimal for large-scale monitoring, reducing computational requirements while maintaining classification performance.
4.5 Inference at scale
A key challenge in promoting sustainable agriculture is understanding adoption patterns of water-saving practices across large regions. The PRANA project trained 150,000 farmers on water-saving techniques, but tracking their implementation was impossible through traditional surveys. Our satellite-based approach enables monitoring at unprecedented scale, providing critical insights for food security policies. To apply our classification framework across Punjab, we first needed to delineate individual agricultural plots. With approximately 3.2 million hectares of rice cultivation and average plot sizes of 0.5-2 hectares, millions of individual fields required identification. We automated this process using the Fields of the World (FTW) algorithm [20], which processes bi-temporal Sentinel-2 imagery to identify field boundaries. Our workflow (Fig. S2) begins with acquiring cloud-free Sentinel-2 imagery from May and October 2024. After preprocessing and dividing the region into manageable tiles, we applied the FTW model to generate field polygons, filtering for size (2,000 $\text{m}^{2}$ to 10 hectares) to match typical smallholder farms. For each identified field, we processed Sentinel-1 time-series data and extracted embeddings from Presto to capture temporal signatures. To address the challenge of sparse training data (1,300 labeled plots within approximately 3 million agricultural plots extracted), we implemented an ensemble approach, selecting the top model from both RF and GB frameworks. These models’ predictions were aggregated by taking the modal value for each rice-growing plot, enhancing robustness and mitigating individual model biases, with the final results in Fig. 1.
Fig. 1 A. shows the spatial distribution of predictions of DSR practicing rice-fields for 2024. It is obvious from visual inspection that there is significant adoption heterogeneity across districts, and even within districts. We validated these results against Punjab government records, which tracked 253,328 acres of DSR cultivation in 2024 through their incentive program [35]. Our district-level predictions showed strong overall correlation with these records (Pearson=0.77, RBO= 0.77444Rank biased overlap (RBO) [36] with p=0.95, standard weighting), though with regional variations (Fig. 1 B). Some discrepancies were observed, with our model underestimating DSR adoption in Fazilka and Sri Muktsar Sahib while overestimating in Sangrur, Patiala, Barnala, Bathinda, and a few other districts, relative to government records. The high RBO indicates a good level of preservation of district rankings, particularly among the highest-ranked districts, complemented by a strong linear correlation between these estimates (Fig. 1 C). These differences may reflect both model limitations and potential inaccuracies in government data, as similar incentive programs in neighboring states have faced verification challenges [37]. For the majority of districts, the model overestimates DSR adoption, except in the top two districts, relative to the provided estimates. The higher RBO relative to Pearson correlation suggests that our model likely captures the relative ordering of districts than absolute values, making it valuable for prioritization exercises and policy applications where identifying relative standings across districts is paramount. Importantly, it enables targeting a sub-district or even at village or plot level. This large-scale application demonstrates how satellite monitoring can provide timely, comprehensive data on sustainable agriculture adoption, enabling more effective policy interventions to address water scarcity and food security challenges.
@@ 5 Limitations
Our approach, while effective, faces several constraints that merit consideration. First, the current Sentinel-1 revisit frequency of 12 days limits our ability to capture high-frequency irrigation variability characteristic of AWD practices. This temporal resolution challenge could be addressed by integrating data from ascending and descending orbits or incorporating forthcoming satellite missions such as Sentinel-1C, which would provide more frequent observations of irrigation cycles.
Second, field boundary extraction using the FTW algorithm is constrained by Sentinel-2’s spatial resolution, which imposes a minimum detectable plot size. This limitation is particularly significant in smallholder farming contexts where field sizes may fall below this detection threshold, potentially excluding smaller agricultural parcels from analysis.
Finally, our models’ performance relies heavily on lag features that capture temporal distinctions between different water management practices. The effectiveness of this approach depends on DSR practices maintaining sufficient temporal separation from other water management methods. Before generalizing these methods to other rice-growing regions globally or even temporally in the same region, additional validation studies must be conducted to determine whether these temporal patterns persist across diverse agroecological contexts and farming calendars.
@@ 6 Conclusion & Future Work
Our study demonstrates that satellite monitoring of rice water management practices is feasible, offering a powerful tool for tracking sustainable agriculture adoption. By classifying practices along separate sowing and irrigation dimensions, we achieved 78% F1-score in distinguishing direct seeded rice from traditional transplanted methods across highly variable planting dates.
This dimensional approach proved critical—DSR classification performed significantly better than attempting to classify all practices simultaneously. The framework successfully scaled to approximately 3 million plots across Punjab, with district-level predictions showing strong correlation (0.79) with government records. Our method requires no prior knowledge of planting dates, making it uniquely suited for real-world agricultural landscapes where cropping schedules vary widely.
While DSR detection performed robustly, monitoring irrigation practices like AWD remains challenging with current satellite revisit frequencies. The upcoming Sentinel-1C mission, expected in 2025, will provide the 6-day temporal resolution needed to better capture irrigation patterns.
This framework directly supports food security initiatives by enabling: 1) evidence-based resource allocation to promote sustainable water management, 2) measurement of intervention impacts at regional and farm scales, 3) targeted support for regions with low adoption rates, and 4) comprehensive monitoring without costly ground surveys
Future work should extend this approach to other major rice-producing regions facing water scarcity. By adapting the framework to diverse agricultural contexts, we can support sustainable intensification globally. Additional opportunities include developing region-specific models, integrating deep learning foundation models like Presto, and creating open data platforms for policymakers and researchers.
As climate change intensifies water scarcity challenges, scalable monitoring of sustainable water management will become increasingly vital for maintaining agricultural productivity and food security. This framework provides a foundation for tracking and promoting practices that allow farmers to produce more food with fewer resources while maintaining regional water security—directly supporting the Zero Hunger Sustainable Development Goal.
@@ Appendix B PRANA Program
We focused our analysis on the Indian state of Punjab, both because of the scale of rice farming, and importance to national and global rice supply, as well as the existing program by The Nature Conservancy, the Promoting Regenerative and No-burn Agriculture (PRANA) project [^5] PRANA website, to promote water-saving rice practices as well and reduce crop residue burning. India is the world’s second biggest producer of rice [38], with the state of Punjab producing 10.42% of the country’s rice in 2024 at 144 million tons[11] representing roughly 3.7% of the world’s rice production capacity. Punjab’s prominence in rice production is heavily supported by irrigation infrastructure and high-yield farming practices but raises concerns about water sustainability due to the heavy reliance on groundwater systems. In Punjab, the adoption of the rice-wheat cropping system has led to a manifold increase in the irrigation water demand, and about 73% of it is met from groundwater [15], with the state experiencing a strong decline groundwater capacity[39] in the last few decades. A large part of this is due to the demand for water for rice cultivation, since the state of Punjab is in a semi-arid region with highly seasonal monsoon precipitation. Thus, unreliable surface water supplies coupled with excessive groundwater exploitation due to free electricity[13] has led to a long-term groundwater decline of 41.6 cm/yr in the state[12]. Han et al. [17] find that rice paddy planting regions increased significantly in northwestern India - one of the regions with the highest conversion from single- to double-cropping of rice.
@@
The PRANA program conducted focused training sessions for ~150,000 farmers on a targeted scale to generate awareness, provide education, agronomic support and monitoring of farmers’ irrigation practices over the years 2023-2024, to enable adoption of both DSR and AWD methods. For both DSR and AWD, potential areas were identified based on the soil type and availability of required machinery. Field demonstrations was conducted in nearly every five villages to showcase the techniques locally. For AWD practice, 27,011 pipes were provided to the farmers pro bono and 26,997 hectares were covered under AWD using these pipes. Apart from these activities, paddy crop guides on agronomic practices were designed and provided to the farmers in local languages to build and strengthen adoption capacity. Fig. 2 shows field photos where these interventions were put into place in 2024.
B.1 Prediction in the Presence of Planting Date Heterogeneity
@@
Previous studies that discriminated DSR from PTR relied on synchronizing SAR backscatter data with cropping calendars, enabling direct modeling of temporal signatures[10, 8, 9]. While effective, this approach is inherently limited in scalability, particularly in regions with unknown or heterogeneous planting dates.
@@
Fig. A1: Planting date distributions across water management practices of interest.
@@
Although planting and harvesting dates were available in our dataset, we intentionally excluded them from the modeling process to enhance the generalizability of our framework. This decision ensures that our approach remains applicable to regions where such information is unavailable. Additionally, farmer interviews revealed the presence of "third crops" (such as lentils, maize, or green manuring crops666a green manure is a crop specifically cultivated to be incorporated into the soil while still green) planted just before the kharif rice season. These crops introduce further variability in SAR backscatter signatures, complicating the detection of water management practices. Due to incomplete data on these practices, third-crop effects were not explicitly modeled but represent an additional source of temporal heterogeneity.
Fig. A2: Examples of hand-crafted features generated from SAR backscatter timeseries for different irrigation field types. Top-left: DSR and CF. Top-right: PTR and AWD. Bottom row: PTR with CF. Left subplots show the VV (blue) and VH (orange) timeseries for every pixel contained within a plot, and the dark blue and dark orange show the mean (with shaded regions denoting the standard deviation) VV and VH, respectively, per plot. Right subplots show the gaussian fit to the VV/VH curve and the radar vegetation index (RVI). The dotted lines in each left-subplot denote the planting date for that plot.

### v2

@@ 4 Results
Dimensional Classification Outperforms Combined Approach
The dimensional approach–classifying sowing (DSR vs. PTR) and irrigation (AWD vs. CF) separately–significantly outperforms the combined three-class formulation, as shown in Table 1. DSR and AWD achieve F1 scores of 0.796 and 0.742, respectively, supporting our hypothesis that practices are more effectively identified along their distinct dimensions, relative to the 3-class formulations. For those same situations, fusing their outcomes reduces the joint accuracy to $\approx 0.6$, confirming that dimensional decomposition avoids error compounding while better reflecting the underlying agronomic processes.
Foundation Model Embeddings Enable Robust Classification
Both Presto and AE yield the highest performance among single feature sources, achieving sowing F1 scores of 0.755 and 0.715, respectively, comparable to the best handcrafted features. Their full value emerges in combination with complementary feature families, particularly Fourier features, where AE+FT attains the best sowing performance (F1=0.798). A similar pattern holds for irrigation, underscoring that foundation model embeddings capture transferable temporal dynamics that enhance orthogonal feature spaces. These results highlight the utility of such embeddings for practitioners, as they are inexpensive to obtain and require no specialized training. Further details are provided in Appendix D.
Lag Features Drive Classification Performance
Lag features encode planting-date offsets between plots. The independent PRANA dataset (Figure A1) shows wide state-level variance but compact practice-specific planting windows, enabling these offsets to act as strong discriminators—though their utility depends on temporal stability across seasons. Practice features capture water-management dynamics intrinsic to each technique. Sowing features (DSR/PTR) are most salient in early growth, while AWD/CF differences emerge later as irrigation frequency diverges (typically every 4–10 days; Figure 4).
To isolate their effects, we trained models on restricted temporal windows (Table 2). Sowing models perform best when both lag and early-season features are available (rows 1, 3, 10–11), confirming that they rely on both signal types and remain robust to sampling frequency (4–7 days). In contrast, irrigation models depend almost entirely on lag information: performance peaks when all features are present (row 11) and collapses once lag cues are reduced (row 8) or sowing phases are absent (rows 5–9). This dependence indicates that AWD detection leverages schedule shifts rather than true irrigation cycles, which Sentinel-1A’s 12-day revisit cannot resolve. Further analyses of feature importance, orbit effects, and errors are provided in Appendix F.
Inference at Scale
A key challenge in promoting sustainable agriculture is understanding adoption patterns of water-saving practices. Despite extensive farmer training, PRANA lacked state-wide visibility due to the high cost of surveys. Our satellite-based approach overcomes this limitation, enabling monitoring of Punjab’s 3.2 million hectares of rice cultivation and providing critical insights for food and water policy. We delineate field boundaries using the FTW algorithm (kerner2024fields) applied to bi-temporal Sentinel-2 RGB-NIR imagery (May & Oct 2024). FTW polygons are filtered to 0.2-10 Ha, matching Punjab’s typical rice-plot size—means: 0.5±0.7 Ha (PRANA internal survey), 0.6±0.5 Ha ( (wang2022unlocking) Punjab subset), and 2.7 Ha (singhEconomicsBasmatiRice2024). Validation of FTW polygons over the PRANA training area shows that 57% of sub-1 Ha plots were detected, with mIoU 0.41 ±0.2 over all plots. FTW often splits large fields ($\approx$1.4 sub-plots per label) and misses 36% of plots, reflecting both its resolution limits and ability to capture smaller sub-plots. We provide further details and qualitative samples of this step in Appendix G and Figure A9.
Given sparse training labels ($\sim$1,400 labeled plots vs. 3 million during inference), we employed an ensemble of top-performing RF and GB models, aggregating predictions via the modal class. Results are summarizes in Figure 1A, showing marked spatial heterogeneity in DSR adoption across and within districts. Our district-level predictions correlate strongly with government records (progressivefarming2025) reporting 253,328 acres under DSR (Spearman’s $\rho$=0.69, RBO=0.77 (webber2010similarity)), despite regional discrepancies. Since our training data included only rice plots, we applied a rice-cultivation mask (han2022annual) from 2021 to exclude non-rice fields. We observe a residual diagonal gradient arises from minor temporal offsets between adjoining S-1 rows, mitigated via radiometric terrain-geocoded correction.
The model generally overestimates DSR area relative to government figures, except in the top three districts. These discrepancies may stem from both model error and noise in the self-reported government data [^3] Data sourced from a DSR incentive program offering Rs. 1500 per acre, pre-verification; c.f. see  (progressivefarming2025) Table 1, which, as seen in neighboring states (Deswal2024DSRFake), may include unverifiable claims. Notably, for Sri Muktsar Sahib and Fazilka, masking introduced large differences ($\sim$40k acres each), likely due to post-2021 expansion in rice cultivation. We therefore report unmasked estimates for these districts (see Table A1 and Appendix G). Further in-field validation will be important to confirm these large-scale adoption estimates. Despite such variations, the high RBO score indicates strong preservation of district rankings, making the system well-suited for policy prioritization by highlighting both lagging and leading districts for targeted action and learning best practices.
@@ 5 Limitations
Our approach, while effective, faces several constraints that merit consideration. First, the current S1 revisit frequency of 12 days limits our ability to capture high-frequency irrigation variability characteristic of AWD practices. This temporal resolution challenge could be addressed by integrating data from ascending and descending orbits or incorporating forthcoming SAR missions.
Second, field boundary extraction using the FTW algorithm is constrained by Sentinel-2’s spatial resolution, which imposes a minimum detectable plot size. This limitation is particularly salient in smallholder farming contexts where field sizes may fall below this detection threshold, potentially excluding smaller agricultural parcels from analysis.
Third, explainability analysis revealed that our models have degrees of reliance on temporal lag features that capture distinctions between water management practices. Additionally, the small dataset size ($\sim$1,400 plots) relative to the large inference dataset underscores the need for careful validation. Although Punjab has relatively low diversity in its agroecological zones, other regions may present greater variation in cropping systems, climate, or practice clustering, affecting generalization. Deployment of such methods would require systematic post-training field validation to verify model predictions against ground truth before operational use, particularly given our reliance on temporal patterns that may shift across agricultural contexts. Therefore, cross-year and cross-region generalization remain future extensions to assess temporal and spatial transferability.
Finally, our analysis depends on external estimates of rice-growing areas. Temporal mismatches between these masks and the study period introduced complications, particularly in regions where rice cultivation has expanded. Future work should explore explicitly classifying non-rice plots to reduce reliance on exteral and outdated maps.
@@ 6 Conclusion & Future Work
Our study demonstrates the feasibility of satellite-based monitoring for rice water management, offering a scalable and data-driven framework for tracking sustainable agriculture adoption. By decoupling classification into sowing (DSR) and irrigation (AWD) dimensions, the approach outperforms combined formulations. Co-designed with agronomy experts, this dimensional structure proved essential: DSR detection notably surpassed joint classification and enabled statewide mapping of DSR adoption across 3 million fields in Punjab, India. Explainability analysis further showed that while AWD classification achieves an relatively high F1 scores, it depends primarily on differential planting schedules rather than irrigation signals, underscoring the importance of interpretability for reliable deployment. AWD performance remains limited by Sentinel-1A’s 12-day revisit, though upcoming missions (Sentinel-1C/D, NISAR) will improve temporal resolution, potentially improving irrigation detection.
Operating without prior knowledge of planting dates, the framework generalizes to real-world systems with heterogeneous cropping calendars. Embeddings from remote sensing foundation models–Presto (tseng2023lightweight) and AlphaEarth (brown2025alphaearth)–were key to this generalization, delivering strong accuracy from limited labeled data. A new class of efficient and highly performant EO foundation models, including OlmoEarth444olmoearth.allenai.org, Panopticon (waldmann2025panopticon) and Galileo (tseng2025galileo), have recently emerged, with precomputed embeddings increasingly available, further extending the potential for scalable, low-cost monitoring of agricultural practices. Finally, future work should examine why AlphaEarth’s year-long embeddings retain strong sensitivity to rice sowing signals despite spanning multiple cropping seasons.
District-level predictions correlate strongly with government records, supporting utility for food and water policy through: (1) evidence-based resource allocation, (2) large-scale impact measurement, (3) identification of low-adoption areas, and (4) statewide monitoring without costly field surveys. At finer scales, plot-level irrigation mapping provides timely, high-resolution data on water management practices, enabling biogeochemical models to more accurately simulate methane emissions and quantify the mitigation benefits of sustainable agriculture. Extending this framework to other rice-growing regions and adapting it to local conditions–crop varieties, irrigation regimes, and management practices–will enhance generalization and policy relevance. As climate change accelerates water scarcity, scalable monitoring of sustainable irrigation will be vital for ensuring food security and advancing the Zero Hunger Sustainable Development Goal.
@@ Appendix B PRANA Program
We focused our analysis on the Indian state of Punjab, both because of the scale of rice farming, and importance to national and global rice supply, as well as the existing program by The Nature Conservancy, the Promoting Regenerative and No-burn Agriculture (PRANA) project, to promote water-saving rice practices as well and reduce crop residue burning. India is the world’s second biggest producer of rice  (AgriculturalProductionStatistics), with the state of Punjab producing 10.42% of the country’s rice in 2024 at 144 million tons (MARKETINTELLIGENCEREPORT2024) representing roughly 3.7% of the world’s rice production capacity. Punjab’s prominence in rice production is heavily supported by irrigation infrastructure and high-yield farming practices but raises concerns about water sustainability due to the heavy reliance on groundwater systems. In Punjab, the adoption of the rice-wheat cropping system has led to a manifold increase in the irrigation water demand, and about 73% of it is met from groundwater (sidhu2021spatio), with the state experiencing a strong decline groundwater capacity (rodell2009satellite) in the last few decades. A large part of this is due to the demand for water for rice cultivation, since the state of Punjab is in a semi-arid region with highly seasonal monsoon precipitation. Thus, unreliable surface water supplies coupled with excessive groundwater exploitation due to free electricity (gupta2023free) has led to a long-term groundwater decline of 41.6 cm/yr in the state (baweja2017groundwater). Han et al.  (han2022annual) find that rice paddy planting regions increased significantly in northwestern India - one of the regions with the highest conversion from single- to double-cropping of rice.
@@
The PRANA program conducted focused training sessions for ~150,000 farmers on a targeted scale to generate awareness, provide education, agronomic support and monitoring of farmers’ irrigation practices over the years 2023-2024, to enable adoption of both DSR and AWD methods. For both DSR and AWD, potential areas were identified based on the soil type and availability of required machinery. Field demonstrations was conducted in nearly every five villages to showcase the techniques locally. For AWD practice, 27,011 pipes were provided to the farmers pro bono and 26,997 hectares were covered under AWD using these pipes. Apart from these activities, paddy crop guides on agronomic practices were designed and provided to the farmers in local languages to build and strengthen adoption capacity. Figure A2 shows field photos where these interventions were put into place in 2024.
Dataset
The dataset used in this study was collected by the authors in collaboration with the PRANA program. Due to the sensitive nature of exact geo-coordinates, and the challenges associated with the traditional method of de-identification (jittering will cause major data fidelity issues), we have chosen to scrub the geo-coordinates from the dataset. Unfortunately, this implies that users wishing to recreate the results will not be able to use the source data (Sentinel-1) to reproduce the steps. Therefore, we provide all the features created for each plot including hand-crafted (HC), Presto (P), Fourtier Transform (FT) and Google AlphaEarth embeddings (AE). These have been packaged with the codebase. Additionally, we provide all code and steps needed for users to generates features for any given geo-referenced polygons.
Prediction in the Presence of Planting Date Heterogeneity
@@
Previous studies that discriminated DSR from PTR relied on synchronizing SAR backscatter data with cropping calendars, enabling direct modeling of temporal signatures (gumma2015mapping; fikriyahDiscriminatingTransplantedDirect2019; villano2019separability). While effective, this approach is inherently limited in scalability, particularly in regions with unknown or heterogeneous planting dates.
@@
Figure A1: Planting date distributions across water management practices of interest.
Figure A2: Water-saving Interventions: A) AWD measuring pipe being installed in a demonstration field. B) AWD pipe after installation. C) Photograph of a sample field practicing DSR taken during the tillering stage; note the lack of flooding in the field.
@@
Although planting and harvesting dates were available in our dataset, we intentionally excluded them from the modeling process to enhance the generalizability of our framework. This decision ensures that our approach remains applicable to regions where such information is unavailable. Additionally, farmer interviews revealed the presence of ”third crops” (such as lentils, maize, or green manuring crops555a green manure is a crop specifically cultivated to be incorporated into the soil while still green) planted just before the kharif rice season. These crops introduce further variability in SAR backscatter signatures, complicating the detection of water management practices. Due to incomplete data on these practices, third-crop effects were not explicitly modeled but represent an additional source of temporal heterogeneity.
Figure A3: Training and validation workflow used in the study. We utilize Copernicus Dataspace and GEE to acquire Sentinel-1 and ERA5 temporal data on a per plot basis, and to subsequently reduce them to plot level statistics. We then fit a spline, resample, generate hand-crafted features and embeddings from Presto over a given time period. The separability of water management classes is modeled using a Random Forest and Gradient Boosted Trees. The inset shows an example of the average S-1 VV and VH timeseries after smoothing and spline fitting, with confidence intervals at each point in time.